A company produces certain types of sophisticated item by three machines. The respective daily production figures are machine A 300 units, Machine B 450 units, and machine C 250 units. Past experience shows that the percentage of defective in the three machines are 0.1,0.2 and 0.7 respectively for machines A B and C. An item is drawn at random from a day's production and is found to be defective. What is the probability that it is not produced by machine C?


1
Expert's answer
2021-03-10T11:00:12-0500

Let "E_1,E_2,E_3" is machine A,B,C


D = defective items


Total production = 300+450+250 = 1000


"P(E_1) = \dfrac{300}{1000}"


"P(E_2) = \dfrac{450}{1000}"


"P(E_3) = \dfrac{250}{1000}"

Now,

"P(\dfrac{D}{E_1}) = 0.001"


"P(\dfrac{D}{E_2}) = 0.002"


"P(\dfrac{D}{E_3}) = 0.007"


Probability = "P(\dfrac{E_3}{D})" = "\dfrac{\dfrac{250}{1000}\times 0.007}{0.001375}"


= "\dfrac{0.000175}{0.001375}"


= "0.127"


P(not produced by machine C ) = 1 - 0.127


= "0.87"